### v2

# Pralekha: Cross-Lingual Document Alignment for Indic Languages

Abstract
Mining parallel document pairs for document-level machine translation (MT) remains challenging due to the limitations of existing Cross-Lingual Document Alignment (CLDA) techniques. Most approaches rely on metadata such as URLs, which is often unavailable in low-resource language settings, while others represent documents using pooled sentence embeddings, which fail to capture fine-grained alignment cues. Moreover, current sentence embedding models have limited context windows, hindering their ability to represent document-level information effectively. To address these challenges for Indic languages, we introduce Pralekha [^1] Pralekha means document in Sanskrit, an ancient Indo-Aryan language spoken in India., a large-scale benchmark for evaluating document-level alignment techniques. It contains over 3 million aligned document pairs across 11 Indic languages and English, of which 1.5 million are English–Indic pairs. Furthermore, we propose Document Alignment Coefficient (DAC), a novel metric for fine-grained document alignment. Unlike pooling-based approaches, DAC aligns documents by matching smaller chunks and computes similarity as the ratio of aligned chunks to the average number of chunks in a pair. Intrinsic evaluation shows that DAC achieves substantial improvements over pooling-based baselines, particularly in noisy scenarios. Extrinsic evaluation further demonstrates that document MT models trained on DAC-aligned pairs consistently outperform those using baseline alignment methods. These results highlight DAC’s effectiveness for parallel document mining. The Pralekha dataset and CLDA evaluation framework will be made publicly available.
Pralekha: Cross-Lingual Document Alignment for Indic Languages
Sanjay Suryanarayanan1   Haiyue Song3 Mohammed Safi Ur Rahman Khan1,4Anoop Kunchukuttan1,2Raj Dabre1,3,4,5thanks: Corresponding author: raj.dabre@cse.iitm.ac.in 1Nilekani Centre at AI4Bharat   2Microsoft 3National Institute of Information and Communications Technology, Japan 4Indian Institute of Technology, Madras   5Indian Institute of Technology, Bombay [Uncaptioned image]  huggingface.co/Pralekha[Uncaptioned image]  github.com/Pralekha
Figure 1: An Overview of the Evaluation Framework for Cross-Lingual Document Alignment. The pink path illustrates the proposed approach leveraging Document Alignment Coefficient (DAC), while the green path represents pooling-based baseline methods.
1 Introduction
The advent of Large Language Models (LLMs) has significantly improved the ability to process long-context information, enabling advancements in document-level tasks such as machine translation (MT). Document-level MT extends beyond sentence-level translation by capturing discourse-level dependencies, coherence, and contextual information across entire documents, leading to more accurate and fluent translations. While there are efforts Ramesh et al. (2022); Siripragada et al. (2020) that provide sentence-level parallel data, large-scale document-level parallel corpora remain scarce Sannigrahi et al. (2023); Pal et al. (2024).
To address this gap, parallel document mining is essential for building high-quality training corpora for document-level MT. However, this progress is hindered by two key challenges: (a) the lack of large-scale document alignment evaluation benchmarks and (b) the limitations of current Cross-Lingual Document Alignment (CLDA) techniques. Reliable evaluation benchmarks are crucial because the effectiveness of an alignment method cannot be improved without accurate performance measurements. As for the limitations of CLDA, most existing embedding models were developed for sentence-level parallel corpus mining Schwenk et al. (2020, 2021); Bañón et al. (2020). As a result, they have limited context windows and can represent only parts of a document Zhu et al. (2024). Document-level embeddings are therefore often obtained by truncating documents or pooling embeddings of independently encoded chunks, both of which can be lossy Sannigrahi et al. (2023). These approaches overlook chunk-level cues, which reduces their ability to capture complete document semantics.
In this paper, we focus on Indic languages, which have abundant sentence-level parallel corpora and benchmarks  Ramesh et al. (2022); Siripragada et al. (2020); Gala et al. (2023) but remain low-resource for CLDA and Document-level MT. We first introduce Pralekha, a large-scale parallel document dataset for evaluating CLDA techniques across 11 Indic languages and English. Pralekha comprises over 3 million document pairs between 11 Indic languages and English, of which 1.5 million are English–Indic pairs. These are high-quality, human-verified aligned documents sourced from reliable domains such as the Indian Press Information Bureau (PIB) and the Mann Ki Baat (MKB) radio program. It serves both as a benchmark for assessing CLDA and as a domain-specific parallel corpus for training document-level MT models for Indic languages. Alongside Pralekha, we propose Document Alignment Coefficient (DAC), a novel metric for CLDA at the fine-granular level. Unlike existing approaches that rely on pooled document-level embeddings, we focus on smaller granular units, called chunks. We first embed and align these chunks, then compute the DAC of a document pair as the ratio of aligned chunks to the average number of chunks in the pair.
In our experiments using Pralekha under realistic noisy document mining settings, we evaluate the proposed DAC approach against existing pooling-based baselines. DAC achieves average precision, recall, and F1 scores of 0.8932, 0.6312, and 0.7372, respectively, showing improvements of 15–20% in precision and 5–10% in F1 score over the baselines. We further benchmark DAC on the CCAligned El-Kishky et al. (2020) dataset, where it attains average precision, recall, and F1 scores of 0.8110, 0.6511, and 0.7203, respectively, outperforming the baselines across all metrics. Having established DAC’s intrinsic effectiveness, we conduct an extrinsic evaluation by training document-level MT models on mined documents using open-source LLMs. The results show that MT models trained on DAC-aligned pairs consistently outperform those trained with baseline alignment methods. The Pralekha dataset and CLDA evaluation framework will be made publicly available.
2 Related Work
Cross-Lingual Sentence Alignment. Early approaches to sentence-level alignment, such as Europarl Koehn (2005), relied on manually curated metadata Abdul-Rauf and Schwenk (2009); Do et al. (2009). Recent advancements in multilingual embeddings have significantly improved large-scale mining of parallel data from unstructured sources. Methods such as LaBSE Feng et al. (2022), LASER Schwenk and Douze (2017), SONAR Duquenne et al. (2023), and hierarchical bilingual retrieval Guo et al. (2019) leverage multilingual sentence representations to enhance cross-lingual alignment accuracy. These techniques have enabled large-scale efforts such as CCMatrix Schwenk et al. (2020), WikiMatrix Schwenk et al. (2021), ParaCrawl Bañón et al. (2020), NLLB Team et al. (2022), and BPCC Gala et al. (2023), which extract and align sentence pairs from massive web crawls.
Cross-Lingual Document Alignment. For document-level alignment, early approaches relied on metadata rather than content, using features such as URL and page structure similarity Resnik and Smith (2003). While multilingual encoders have improved sentence-level parallel data extraction, document-level alignment remains challenging due to the limited context windows of existing sentence embedding models Zhu et al. (2024). In this work, we explore a chunk-based method that optimally leverages fine-granular embeddings for document-level alignment.
Indic Language Corpora. Despite India’s large population (1.43 billion), its languages remain low-resource, particularly for high-quality document-level data. Significant efforts have been made to build monolingual corpora, such as IndicNLP Suite Kakwani et al. (2020) and Sangraha Khan et al. (2024), as well as parallel corpora and benchmarks like IN22 and BPCC Gala et al. (2023). However, no publicly available cross-lingual parallel document datasets exist for Indic languages Sannigrahi et al. (2023). The only work referencing parallel documents uses them solely for mining sentence pairs, without releasing the document-level data Siripragada et al. (2020). This underscores the urgent need to construct parallel document corpora for cross-lingual document alignment and document-level machine translation, a gap that our work aims to fill.
3 Pralekha
Figure 2: Heat-map of alignable document pairs for each language pair in Pralekha. Darker cells indicate higher alignment counts.
We present Pralekha a large-scale parallel document dataset for evaluating Cross-Lingual Document Alignment (CLDA) techniques. Pralekha comprises over 3 million document pairs between 12 languages - Bengali (ben), Gujarati (guj), Hindi (hin), Kannada (kan), Malayalam (mal), Marathi (mar), Odia (ori), Punjabi (pan), Tamil (tam), Telugu (tel), Urdu (urd) and English (eng) containing a mixture of high, medium and low resource languages. Figure 2 shows a heatmap of the distribution of alignable document pairs across language combinations. Among these, 1.5 million document pairs are English–Indic.
Figure 3: Average document length in Pralekha for each language, measured in terms of number of sentences per document.
Figure 3 presents statistics on the average number of sentences per document for each language in Pralekha. The average sentence count ranges between 20 and 30, indicating that the documents are relatively long, making our benchmark challenging for CLDA.
Pralekha covers two broad domains - News Bulletins and Podcast Scripts, therefore containing a mixture of written as well as spoken form of data. Following the methodology of Sangraha Khan et al. (2024), we custom-scraped data from the Indian Press Information Bureau (PIB) [^2] https://pib.gov.in website, aligning documents by matching bulletin IDs interlinking bulletins across languages. These documents are manually written and hence are of high quality. For podcast scripts, we employed the approach used by Khan et al. (2024) and Siripragada et al. (2020) to collect transcripts from Mann Ki Baat, [^3] https://www.pmindia.gov.in/en/mann-ki-baat a radio program hosted by the Indian Prime Minister. This program is typically spoken in Hindi, then manually transcribed and translated into various other Indian languages. Sometimes, the website’s meta-data incorrectly marks the document’s languages, which we fix with a simple combination of IndicLID Madhani et al. (2023) and script-specific Unicode ranges.
All the data is human-written/verified. This dataset serves both as a benchmark for assessing document alignment techniques and as a domain-specific parallel corpus for training document-level MT models in Indic Languages.
4 Cross-Lingual Document Alignment
Given two document collections, $D_{1}$ and $D_{2}$, where $D_{1}$ consists of documents written in lang1 and $D_{2}$ consists of documents in lang2, the objective of Cross-Lingual Document Alignment (CLDA) is to determine a set of paired documents, $P=\{(d_{1},d_{2})\mid d_{1}\in D_{1},d_{2}\in D_{2}\}$, such that each pair $(d_{1},d_{2})$ exhibits a high degree of semantic similarity, with $d_{1}$ and $d_{2}$ being direct translations of each other.
We next describe the evaluation framework used to benchmark various CLDA techniques. As illustrated in Figure 1, the framework includes our proposed approach, which leverages the Document Alignment Coefficient (DAC), along with the pooling-based baseline methods. The following subsections discuss each component of the framework in detail.
4.1 Granularity
The granularity level $G$ defines the unit of text at which alignment is performed and is a key parameter in our study. We evaluate document alignment at three levels of granularity: $G=1$ (sentence-level, as in Schwenk et al. (2020)), $G=2,4,8$ (chunk-level, where each chunk consists of 2, 4, or 8 sentences, respectively), and $G=|D|$ (document-level, as in El-Kishky et al. (2020); El-Kishky and Guzmán (2020)).
At one extreme, $G=1$ produces too many embeddings to align. This makes the process inefficient and provides limited context for semantic comparison. At the other extreme, $G=|D|$ leads to overly coarse representations and information loss. It also creates challenges in computing a single document-level embedding because embedding models have a limited context window Zhu et al. (2024). Both extremes have limitations; therefore, we consider intermediate values of $G$ that strike a balance between efficiency and information retention. A larger granularity helps preserve local coherence and capture broader document-level information, addressing the shortcomings of traditional sentence-based approaches Xie et al. (2024).
Pooling Strategies for Baselines. We benchmark several pooling strategies used in prior work El-Kishky et al. (2020); El-Kishky and Guzmán (2020) to construct document embeddings ($G=|D|$) by aggregating finer-grained unit embeddings.
1. Mean Pooling (MP): Computes a single embedding by averaging all unit embeddings.
2. Length Pooling (LP): Adjusts unit embeddings based on length, assigning greater importance to longer, more informative segments.
3. Inverse Document Frequency (IDF): It weighs terms according to their rarity, prioritizing informative words while down-weighting common ones to improve alignment precision.
4. Length-Inverse Document Frequency (LIDF): Combines length-based weighting and IDF, emphasizing longer, content-rich segments while reducing the influence of repetitive or generic terms, providing a balanced approach to alignment.
4.2 Embedding Model
To compute multilingual embeddings we consider LaBSE [Uncaptioned image] Feng et al. (2022) and SONAR [Uncaptioned image] Duquenne et al. (2023). LaBSE is designed for cross-lingual sentence retrieval, leveraging both parallel and monolingual data to generate language-agnostic embeddings. SONAR extends this approach by training on a more diverse set of languages, leading to improved representation quality for both high-resource and low-resource languages. Computing Embeddings. After segmenting each document into units based on the chosen granularity $G$, we obtain representations for each segment by encoding the entire unit with the embedding model. Because the values of $G$ considered in our study are relatively small, each segment fits within the context window of the embedding models used.
4.3 Alignment Algorithm
[…]
Algorithm 1 Alignment Algorithm
1:Input: Embeddings $\mathbf{X}=\{x_{i}\}$ and $\mathbf{Y}=\{y_{j}\}$, $k$ nearest neighbors
2:Output: $\mathcal{U}=\arg\max_{\mathcal{U}^{\prime}}\sum_{(x,y)\in\mathcal{U}^{\prime}}M(x,y)$
3:FAISS Indexing: Construct FAISS index for efficient similarity search
[…]
$\texttt{NN}_{k}(x)=\{y_{1},y_{2},\dots,y_{k}\}\subset\mathbf{Y}\text{ for each }x\in\mathbf{X}$
$\texttt{NN}_{k}(y)=\{x_{1},x_{2},\dots,x_{k}\}\subset\mathbf{X}\text{ for each }y\in\mathbf{Y}$
5:Compute Margin Score:
6:for each $(x,y)$ in $\mathbf{X}\times\mathbf{Y}$ where $y\in\texttt{NN}_{k}(x)$ or $x\in\texttt{NN}_{k}(y)$ do
$A_{x}=\frac{1}{k}\sum_{z\in\texttt{NN}_{k}(x)}\cos(x,z)$
$A_{y}=\frac{1}{k}\sum_{z\in\texttt{NN}_{k}(y)}\cos(y,z)$
$M(x,y)=\frac{\cos(x,y)}{0.5\times(A_{x}+A_{y})}$
7:end for
8:Max Strategy: Select pairs to maximize overall margin score
$\mathcal{U}=\left\{(x,y)\mid y\in\texttt{NN}_{k}(x)\right\}$
$\cup\left\{(x,y)\mid x\in\texttt{NN}_{k}(y)\right\}$
$\mathcal{U}_{\text{sorted}}=\text{Sort}(\mathcal{U})\text{ by }M(x,y)\text{ in descending order}$
9:for each $(x,y)$ in $\mathcal{U}_{\text{sorted}}$ do
10:  if both $x$ and $y$ are not in aligned pairs then
11:   Add $(x,y)$ to aligned pairs
12:  end if
13:end for
14:Return: $\mathcal{U}$, aligned pairs that maximize the overall margin score
For pooled document embeddings at $G=|D|$, we directly apply the alignment algorithm on these embeddings to obtain aligned document pairs. We extend this idea by introducing a novel Document Alignment Coefficient (DAC) metric for mining parallel documents, which leverages finer-grained units for document alignment. The motivation behind DAC is to avoid relying solely on document-level embeddings, which are typically formed by pooling unit embeddings and may fail to capture detailed semantic information. Instead, we treat each document as a sequence of chunks, where each chunk is a contiguous set of sentences and serves as the basic unit of alignment. Our hypothesis is that embeddings at finer granularity capture semantic nuances more effectively. After performing alignment at the chunk level, the resulting information is aggregated to identify aligned document pairs. The following subsection provides a detailed description of the DAC approach.
4.4 Document Alignment Coefficient
Unlike baseline approaches that operate directly on pooled document embeddings, DAC performs chunk-level alignment first and then derives document pairs using a computed score based on these chunk alignments, thereby providing finer-grained semantic signals that improve alignment precision.
Computing DAC for a Document Pair.
For a given source–target document pair, the DAC is computed as follows:
$\text{DAC}=\frac{2\times N_{\text{aligned}}}{N_{\text{src}}+N_{\text{tgt}}}$ (1)
Here, $N_{\text{src}}$ and $N_{\text{tgt}}$ represent the total number of chunks in the source and target documents, respectively, and $N_{\text{aligned}}$ denotes the number of aligned chunks between the two documents.
The algorithm first aligns the constituent chunks within each document pair and then uses these fine-grained alignments to compute the overall degree of alignment at the document level.
DAC Threshold.
The DAC produces a normalized alignment score between 0 and 1, where higher values indicate stronger alignment, and a threshold on this score can be set to control the final alignment quality. A lower DAC threshold allows a balance between quality and quantity, accepting some misalignments while increasing coverage. On the other hand, a higher threshold prioritizes precision, ensuring fewer incorrect alignments at the cost of reduced yield. The threshold can be adjusted based on the specific requirements of the aligned dataset and its intended downstream task. A detailed analysis of various DAC Thresholds is provided in Figure 7 and Table 3 in Appendix A
Language ben guj hin kan mal mar ori pan tam tel urd eng
Alignable 95K 67K 204K 61K 67K 134K 46K 108K 149K 109K 220K 298K
Unalignable 47K 34K 102K 31K 34K 67K 23K 54K 75K 55K 110K 149K
Total 142K 101K 306K 92K 101K 201K 69K 162K 224K 164K 330K 447K
Table 1: Per language statistics of unique documents divided into two groups: alignable and unalignable. Alignable documents come from the English-Indic part of Pralekha and the Unalignable documents come from Sangraha Unverified, representing noise for a realistic evaluation setting.
5 Experiments
In this section, we present the experimental setup and describe the two types of evaluations conducted on Pralekha: intrinsic evaluation, which measures the effectiveness of document alignment techniques, and extrinsic evaluation, which assesses their impact on downstream tasks. Unless stated otherwise, all references to Pralekha hereafter refer to its English–Indic subset.
5.1 Intrinsic Evaluation
For intrinsic evaluation, we assess how well DAC performs compared to existing pooling-based baselines for document alignment.
Data Setting :
Pralekha contains only parallel (alignable) documents, so aligning only these would not reflect a realistic and challenging scenario. To simulate real-world multilingual corpora setting, we sample random (unalignable) documents from Sangraha Unverified Khan et al. (2024) that cannot be aligned with those in Pralekha and inject them as noise into the dataset. We select a number of unalignable documents equal to 50% of the documents in Pralekha, resulting in a 1:2 ratio of unalignable to alignable documents during alignment. Table 1 reports the counts of unique alignable and unalignable documents for each language used in our experiments. Additionally, we compare the best-performing baseline with our proposed DAC approach on the CCAligned dataset El-Kishky et al. (2020), using a similar data setting.
Alignment Settings :
[…]
Evaluation Metrics :
To assess the intrinsic performance of various document alignment methods, we evaluate three key metrics: precision, recall, and F1 score. While a high F1 score is generally desirable as it balances precision and recall, the ideal metric depends on the specific downstream task.
5.2 Extrinsic Evaluation
To assess the extrinsic performance of document alignment techniques, we evaluate their effectiveness in document-level MT tasks.
Data Setting :
We focus on translation between English and eight Indic languages: Bengali, Gujarati, Hindi, Kannada, Malayalam, Marathi, Odia, and Tamil. From Pralekha, we randomly sample 1,000 document pairs each for the development and test sets. After removing these pairs, we perform document alignment on the remaining data as described in Section 5.1. We obtain two sets of aligned documents: one using the best-performing baseline and the other using our proposed DAC approach. From each set, we sample 10,000 aligned pairs to train two separate document-level MT models using open-source LLMs. Extrinsic evaluation is then conducted on the sampled test set.
Implementation and Training :
We fine-tune two open-source LLMs: Llama-3.2-1B666huggingface.co/meta-llama/Llama-3.2-1B [Uncaptioned image] and Sarvam-1777huggingface.co/sarvamai/sarvam-1 [Uncaptioned image]. LLaMA-3.2-1B is a 1 billion parameter multilingual model released by Meta AI, while Sarvam-1 is a 2 billion parameter model developed by Sarvam AI, optimized for Indic languages. We used open-instruct [^8] https://github.com/allenai/open-instruct for full fine-tuning using the prompt-completion template. We used default hyperparameters recommended in open-instruct, most notably: learning rates of $5\times 10^{-5}$, maximum context (prompt+completion) lengths of 4096, Adam optimizer and no weight decay. We trained only bilingual models, a single model for English $\leftrightarrow$ X translation, for a total of 16 models (8 languages, 2 alignment approaches). Our models were trained for 1 epoch. We did greedy decoding with a maximum of 4,096 generated tokens.
Metrics Methods LaBSE[Uncaptioned image] SONAR[Uncaptioned image]
G = 1 G = 2 G = 4 G = 8 G = 1 G = 2 G = 4 G = 8
Precision MP 0.7585 0.7432 0.7232 0.7018 0.7803 0.7643 0.7541 0.7632
LP 0.7820 0.7709 0.7614 0.7464 0.7943 0.7843 0.7891 0.8111
IDF 0.7610 0.7463 0.7290 0.7052 0.7821 0.7669 0.7596 0.7661
LIDF 0.7831 0.7716 0.7619 0.7466 0.7950 0.7849 0.7894 0.8112
DAC 0.9152 0.9007 0.8936 0.8912 0.9106 0.8870 0.8742 0.8732
Recall MP 0.6310 0.6057 0.5653 0.5599 0.6613 0.6349 0.5696 0.5414
LP 0.6824 0.6807 0.7019 0.6959 0.6924 0.6834 0.6688 0.6630
IDF 0.6437 0.6192 0.5962 0.5738 0.6702 0.6471 0.5991 0.5544
LIDF 0.6908 0.6844 0.7030 0.6971 0.6953 0.6864 0.6700 0.6632
DAC 0.6588 0.6411 0.6471 0.6407 0.6484 0.6108 0.6005 0.6020
F1 Score MP 0.6856 0.6639 0.6291 0.6157 0.7133 0.6904 0.6410 0.6237
LP 0.7252 0.7199 0.7273 0.7170 0.7372 0.7276 0.7213 0.7266
IDF 0.6940 0.6733 0.6519 0.6266 0.7193 0.6989 0.6647 0.6353
LIDF 0.7303 0.7223 0.7281 0.7178 0.7392 0.7297 0.7221 0.7268
DAC 0.7635 0.7463 0.7480 0.7428 0.7550 0.7210 0.7098 0.7109
Table 2: Average Precision, Recall, and F1 scores from the Intrinsic Evaluation of CLDA techniques on Pralekha, computed over 11 Indic languages. Results are shown for LaBSE and SONAR embeddings at granularities $G=1,2,4,8$. We compare pooling-based baselines (MP, LP, IDF, LIDF) with our proposed DAC approach. Bold values indicate the best score for each granularity. Appendix A provides further analysis on the intrinsic trends.
Evaluation Metrics : To evaluate document-level MT performance, we use DocCOMET [^9] https://github.com/Unbabel/COMET, an extension of COMET that incorporates document-level context for improved quality assessment. Specifically, we employ the Unbabel/wmt22-comet-da model from the WMT 2022 Metrics Shared Task, a reference-based COMET model designed for document-level MT evaluation. Alongside COMET, we also report ChrF [^10] https://github.com/mjpost/sacreBLEU#chrf--chrf, a character-level F-score metric based on n-gram overlaps.
6 Results and Discussion
We present results demonstrating the effectiveness of DAC through intrinsic and extrinsic evaluations.
6.1 Intrinsic Performance of DAC
Table 2 reports average precision, recall, and F1 scores across granularities and embedding models.
DAC vs. Baselines:
DAC consistently achieves the highest precision across all settings, with gains of 15–20% over pooling-based baselines. While its recall is slightly lower, DAC attains competitive or superior F1 scores, indicating that it produces more precise alignments, a desirable property for high-quality MT training data, where quality is more important than quantity Ranathunga et al. (2024); Zhou et al. (2023). Among the baselines, LIDF is the strongest competitor, performing closest to DAC due to its effective combination of length normalization and term weighting. LP performs slightly worse, whereas MP and IDF-based pooling lag considerably behind. Additionally, in experiments conducted on the CCAligned dataset El-Kishky et al. (2020), DAC outperforms LIDF across all metrics; see Appendix B for more details.
[…]
Effect of Granularity and Embedding Model:
Performance decreases slightly as $G$ increases, with the largest drop from $G=1$ to $G=2$; differences between $G=4$ and $G=8$ are minimal. This trade-off is expected, as higher granularities reduce the number of chunks to align, enabling faster mining at a small cost to quality. Among embedding models, SONAR performs better with pooling-based methods, whereas LaBSE achieves higher performance with DAC. This aligns with their training setups: SONAR is trained with mean-pooled embeddings, while LaBSE uses CLS representations, which are more effective for fine-grained alignment.
In conclusion, the optimal configuration for document alignment is LaBSE with DAC at granularities $G=1$, $G=2$, and $G=4$, followed closely by SONAR with LIDF at the same granularities. See Appendix A for a more detailed analysis of intrinsic performance trends.
Based on these findings, we use the selected methods for extrinsic evaluation.
6.2 Extrinsic Performance of DAC
Table 5 reports COMET scores for extrinsic evaluation, comparing DAC with LIDF, the strongest baseline from Section 6.1. DAC consistently outperforms LIDF across all granularities, translation directions, and fine-tuned models, showing that its higher-precision alignments produce better document pairs and improve translation quality. Detailed per-language COMET scores are also provided in Table 5 in the Appendix.
Differences in MT performance across granularities are small, consistent with intrinsic trends. The strong results for $G=2$ and $G=4$ indicate that chunk-based alignment can reduce computation while maintaining quality.
[…]
Overall, performance differences are modest but consistent. Extrinsic results show that DAC-aligned documents at $G=2$ and $G=4$ achieve strong translation quality, while intrinsic performance does not always predict downstream gains.
Appendix C provides a more detailed analysis of extrinsic performance trends. Additionally, we report ChrF scores in Table 6, which follow trends similar to the COMET results.
7 Conclusion
Mining parallel document pairs for document-level MT is challenging because existing CLDA methods often rely on metadata or pooled sentence embeddings due to the limited context windows of embedding models. To address these challenges for Indic languages, we introduced Pralekha, a large-scale benchmark containing over 3 million aligned document pairs across 11 Indic languages and English. We also proposed the Document Alignment Coefficient (DAC), a fine-grained alignment metric that leverages chunk-level matches instead of traditional pooling-based representations. Intrinsic evaluation in realistic noisy settings shows that DAC outperforms existing baselines, achieving higher precision and F1 scores. Extrinsic evaluation further demonstrates that document-level MT models trained on DAC-aligned pairs consistently outperform those trained on baseline alignments, validating DAC’s effectiveness for parallel document mining. By releasing the Pralekha dataset and CLDA evaluation framework, we aim to enable further research on developing and evaluating document-level MT models for Indic languages.
Limitations
While Pralekha and the proposed Document Alignment Coefficient (DAC) advance research in cross-lingual document alignment (CLDA), several limitations remain. Our dataset covers 11 Indic languages and English, so the findings may not generalize to other language families with different linguistic properties, such as complex morphology or divergent word order. The documents in Pralekha are mainly drawn from structured, high-quality sources (PIB and MKB). As a result, the benchmark is less representative of noisy web data that contains informal language, OCR errors, or code-switching. In addition, the benchmark primarily includes documents of moderate length. Performance on much shorter or longer documents, which are common in other domains, remains untested. The DAC metric, though more accurate than pooling-based methods, is computationally expensive because it performs chunk-level alignment. This cost may limit its use for very large corpora or real-time applications. DAC may also be less robust when translations are semantically equivalent but highly paraphrased. Finally, our extrinsic evaluation does not fully establish how improvements in alignment quality affect document-level MT performance, since current MT metrics capture discourse-level phenomena only to a limited extent. Future work should expand CLDA evaluation to more diverse language families, noisier domains, and a wider range of document lengths. It should also explore alignment strategies that are more robust to paraphrasing, improve computational efficiency, and develop document-level MT metrics that better correlate with alignment quality.
Ethics
The code and datasets created in this work will be made available under permissible licenses. Generative AI systems were only used for assistance purely with the language of the paper, eg: paraphrasing, spell-check, polishing the author’s original content and for writing boiler plate code.
Appendix
Appendix A Intrinsic Performance Trends in CLDA
This section analyzes the Precision, Recall, and F1 score trends of various Cross-Lingual Document Alignment (CLDA) methods based on their intrinsic performance on Pralekha, as shown in Table 2.
A.1 Impact of Granularities
Figure 5: Impact of Granularity on the Intrinsic Performance of CLDA techniques on Pralekha.
As shown in Figure 5, increasing granularity is associated with a gradual reduction in precision, recall, and F1 score. The decline is most pronounced for recall, suggesting that fewer relevant matches are retrieved as segmentation becomes finer. Precision, however, remains relatively stable across granularities. These trends indicate that while finer segmentation can provide more detailed units for alignment, it may also reduce the amount of contextual information available to embedding models, which are typically optimized for shorter text spans.
A.2 Impact of Embedding Models
Figure 6: Impact of Embedding Models on the Intrinsic Performance of CLDA techniques on Pralekha.
Figure 6 shows how increasing granularity affects the performance of both LaBSE and SONAR across precision, recall, and F1 score. SONAR tends to maintain slightly higher precision, indicating more accurate alignments, whereas LaBSE achieves higher recall, retrieving a larger number of matches. At finer granularities, both models show a gradual decrease in overall performance, with SONAR being relatively more stable in precision. LaBSE achieves a higher F1 score overall, reflecting a more balanced trade-off between precision and recall. These results suggest that SONAR may be better suited for scenarios where precision is prioritized, while LaBSE can be advantageous for broader retrieval needs.
A.3 Impact of DAC Threshold
Figure 7: Impact of DAC Threshold on the Intrinsic Performance of CLDA techniques on Pralekha.
Figure 7 illustrates the trends of precision, recall, and F1 score across different DAC thresholds, highlighting the trade-offs between precision and recall. Increasing the DAC threshold leads to higher precision but lower recall and F1 scores in all settings. We adopt a threshold of 0.1 for both intrinsic and extrinsic evaluations, as it provides a better balance between quality and yield. Table 3 compares the intrinsic performance of DAC at thresholds 0.1 and 0.5.
Metrics Methods LaBSE[Uncaptioned image] SONAR[Uncaptioned image]
G = 1 G = 2 G = 4 G = 8 G = 1 G = 2 G = 4 G = 8
Precision DAC (0.1) 0.9152 0.9007 0.8936 0.8912 0.9106 0.8870 0.8742 0.8732
DAC (0.5) 0.9555 0.9617 0.9570 0.9404 0.9546 0.9600 0.9480 0.9307
Recall DAC (0.1) 0.6588 0.6411 0.6471 0.6407 0.6484 0.6108 0.6005 0.6020
DAC (0.5) 0.2930 0.2568 0.3104 0.3960 0.2530 0.2005 0.2542 0.3297
F1 Score DAC (0.1) 0.7635 0.7463 0.7480 0.7428 0.7550 0.7210 0.7098 0.7109
DAC (0.5) 0.4368 0.3915 0.4564 0.5486 0.3896 0.3229 0.3931 0.4814
Table 3: Comparison of DAC threshold performance (0.1 and 0.5) in the intrinsic evaluation on Pralekha. Bold values indicate the best score for each granularity.
Figure 8: Intrinsic evaluation performance of DAC, LP, and LIDF on Pralekha.
A.4 DAC vs. Baseline Methods
We compare our DAC approach with the two strongest pooling-based baselines, Length Pooling (LP) and Length-IDF (LIDF), as shown in Table 2. Figure 8 highlights that DAC consistently delivers the highest precision across all granularities for both LaBSE and SONAR. Although LP and LIDF yield slightly higher recall, DAC achieves a better balance between precision and recall, reflected in its F1 scores. Moreover, DAC shows minimal degradation in precision as granularity increases, demonstrating its robustness compared to the baselines. These trends confirm that DAC more effectively leverages contextual information, making it the strongest overall alignment method.
[…]
Appendix B Intrinsic Evaluation on CCAligned
Language
Method
Precision Recall F1
G = 1 G = 2 G = 4 G = 1 G = 2 G = 4 G = 1 G = 2 G = 4
Assamese LIDF 0.4596 0.4842 0.4743 0.3511 0.3862 0.3781 0.3981 0.4297 0.4208
DAC 0.5978 0.5792 0.5635 0.3966 0.3840 0.3497 0.4768 0.4618 0.4316
Bengali LIDF 0.7569 0.7870 0.7785 0.6791 0.7274 0.7153 0.7159 0.7560 0.7456
DAC 0.8330 0.8317 0.8358 0.6848 0.7125 0.6961 0.7517 0.7675 0.7596
Gujarati LIDF 0.8206 0.8248 0.8138 0.7660 0.7768 0.7646 0.7924 0.8001 0.7884
DAC 0.8967 0.8768 0.8559 0.7833 0.7671 0.7375 0.8390 0.8183 0.7923
Kannada LIDF 0.7042 0.7079 0.7001 0.6399 0.6470 0.6415 0.6705 0.6761 0.6695
DAC 0.8240 0.8043 0.7825 0.6625 0.6410 0.6069 0.7345 0.7134 0.6836
Malayalam LIDF 0.7196 0.7247 0.7090 0.6350 0.6308 0.6113 0.6747 0.6745 0.6565
DAC 0.8417 0.8116 0.7816 0.7032 0.6717 0.6172 0.7662 0.7350 0.6898
Marathi LIDF 0.7741 0.7787 0.7738 0.7121 0.7194 0.7110 0.7418 0.7478 0.7411
DAC 0.8591 0.8368 0.8122 0.7359 0.7109 0.6748 0.7928 0.7687 0.7372
Odia LIDF 0.5793 0.5739 0.5566 0.4014 0.4036 0.3886 0.4742 0.4739 0.4576
DAC 0.7322 0.7326 0.7057 0.4262 0.4270 0.3991 0.5388 0.5395 0.5099
Tamil LIDF 0.7115 0.7079 0.6976 0.6493 0.6413 0.6332 0.6789 0.6729 0.6638
DAC 0.8265 0.8034 0.7775 0.6628 0.6404 0.6006 0.7356 0.7127 0.6777
Telugu LIDF 0.7416 0.7519 0.7517 0.6758 0.6905 0.6877 0.7072 0.7199 0.7183
DAC 0.8359 0.8145 0.7962 0.7081 0.6852 0.6515 0.7667 0.7443 0.7166
Table 4: Precision, Recall, and F1 scores from Intrinsic Evaluation on CCAligned at varying granularities ($G$). We compare the best-performing pooling-based baseline from Table 2 (LIDF) with our proposed approach (DAC). Bold values indicate the best score for each granularity.
While the primary intrinsic evaluation was conducted on the Pralekha dataset, we also perform an additional benchmark on CCAligned El-Kishky et al. (2020), a large-scale web-mined parallel corpus. This experiment evaluates whether the performance gains of DAC generalize beyond the structured and high-quality domains of Pralekha. We follow the same data settings described in Section 5.1. For comparison, we use the strongest baseline from the main experiments: the LIDF pooling method with SONAR embeddings, evaluated at granularities $G=1,2,4$ and the DAC approach with LaBSE embeddings at the same granularities.
Table 4 presents the results, showing that DAC consistently outperforms LIDF across all metrics and settings. Its effectiveness is therefore not limited to Pralekha but also extends to large, noisy web-mined corpora, confirming DAC as a robust alignment method across diverse domains.
Appendix C Extrinsic Performance Trends in CLDA
Alongside COMET, we report ChrF scores. COMET uses neural models to capture contextual information, whereas ChrF measures surface-level lexical similarity, providing a complementary perspective on translation quality.
Figure 9 shows COMET and ChrF scores across granularities ($G=1,2,4$) for the Llama 3.2-1B and Sarvam-1 models. Sarvam-1 consistently outperforms Llama 3.2-1B across all granularities, and ChrF follows trends similar to COMET as discussed in Section 6.2. DAC-aligned documents at $G=2$ and $G=4$ slightly outperform LIDF in both metrics, confirming DAC’s effectiveness for document alignment. These results also highlight that strong intrinsic performance does not always translate to gains in downstream tasks.
English $\rightarrow$ Indic Indic $\rightarrow$ English
Language Method Llama-3.2-1B [Uncaptioned image] Sarvam-1-2B [Uncaptioned image] Llama-3.2-1B [Uncaptioned image] Sarvam-1-2B [Uncaptioned image]
G = 1 G = 2 G = 4 G = 1 G = 2 G = 4 G = 1 G = 2 G = 4 G = 1 G = 2 G = 4
Bengali LIDF 0.7242 0.7277 0.7112 0.8748 0.8738 0.8711 0.8131 0.8079 0.8109 0.8283 0.8212 0.8235
DAC 0.7693 0.7694 0.7548 0.8764 0.8778 0.8755 0.8270 0.8168 0.8217 0.8331 0.8307 0.8296
Gujarati LIDF 0.7315 0.7036 0.6771 0.8907 0.8897 0.8925 0.8467 0.8484 0.8447 0.8622 0.8627 0.8603
DAC 0.7326 0.7458 0.7379 0.8944 0.8905 0.8966 0.8485 0.8541 0.8541 0.8640 0.8702 0.8677
Hindi LIDF 0.7874 0.7823 0.7849 0.8334 0.8336 0.8324 0.8525 0.8561 0.8552 0.8541 0.8520 0.8552
DAC 0.7894 0.7848 0.7908 0.8357 0.8338 0.8366 0.8584 0.8562 0.8553 0.8558 0.8531 0.8554
Kannada LIDF 0.6340 0.6202 0.6317 0.8707 0.8699 0.8702 0.8257 0.8272 0.8231 0.8387 0.8392 0.8435
DAC 0.6643 0.6417 0.6619 0.8713 0.8771 0.8717 0.8334 0.8290 0.8284 0.8437 0.8446 0.8450
Malayalam LIDF 0.6829 0.6793 0.6598 0.8794 0.8815 0.8809 0.8229 0.8218 0.8279 0.8432 0.8419 0.8397
DAC 0.7139 0.7139 0.6958 0.8848 0.8825 0.8831 0.8315 0.8340 0.8342 0.8496 0.8442 0.8453
Marathi LIDF 0.6552 0.6553 0.6479 0.7424 0.7403 0.7407 0.8241 0.8163 0.8223 0.8501 0.8395 0.8412
DAC 0.6828 0.6701 0.6655 0.7474 0.7528 0.7442 0.8289 0.8307 0.8308 0.8578 0.8529 0.8547
Odia LIDF 0.5142 0.5328 0.5394 0.8761 0.8752 0.8767 0.8241 0.8163 0.8223 0.8501 0.8395 0.8412
DAC 0.6391 0.6233 0.5910 0.8778 0.8754 0.8766 0.8362 0.8427 0.8295 0.8509 0.8545 0.8508
Tamil LIDF 0.7112 0.7177 0.7131 0.8927 0.8929 0.8934 0.7846 0.7917 0.7892 0.8115 0.8089 0.8082
DAC 0.7559 0.7525 0.7518 0.8990 0.9024 0.8996 0.7979 0.8056 0.8029 0.8180 0.8186 0.8167
[…]
English $\rightarrow$ Indic Indic $\rightarrow$ English
Language Method Llama-3.2-1B [Uncaptioned image] Sarvam-1-2B [Uncaptioned image] Llama-3.2-1B [Uncaptioned image] Sarvam-1-2B [Uncaptioned image]
G = 1 G = 2 G = 4 G = 1 G = 2 G = 4 G = 1 G = 2 G = 4 G = 1 G = 2 G = 4
Bengali LIDF 34.17 34.22 33.52 48.57 48.79 48.78 57.32 55.40 55.98 68.34 68.75 68.68
DAC 39.25 38.43 37.47 50.62 50.06 50.90 58.59 55.71 57.20 69.95 69.68 69.74
Gujarati LIDF 35.11 32.84 31.60 64.06 63.48 63.13 36.12 38.54 36.29 64.31 66.37 65.81
DAC 36.12 38.54 36.29 64.31 66.37 65.81 40.75 41.80 40.95 67.80 68.21 67.92
Hindi LIDF 57.32 55.40 55.98 68.34 67.75 67.68 37.20 36.90 36.50 49.80 49.60 49.90
DAC 58.59 55.71 57.20 69.95 68.68 68.74 42.00 41.60 41.20 52.40 52.10 52.50
Kannada LIDF 29.71 29.01 29.47 62.91 63.42 62.81 33.20 32.80 32.40 45.80 45.60 46.00
DAC 32.70 31.23 32.15 64.84 64.66 64.82 37.80 37.40 37.00 48.60 48.30 48.70
Marathi LIDF 40.70 39.79 40.28 55.86 55.61 54.76 35.50 35.10 34.60 47.50 47.20 47.70
DAC 45.53 43.55 42.94 60.02 58.30 57.17 40.00 39.50 39.00 50.40 50.10 50.60
Malayalam LIDF 33.65 33.65 31.94 57.65 58.08 58.60 33.90 33.50 33.10 45.90 45.60 46.10
DAC 37.03 36.59 35.40 62.44 59.86 60.35 38.20 37.80 37.40 48.50 48.20 48.60
Odia LIDF 23.00 23.77 24.42 57.65 56.15 56.85 33.90 33.50 33.10 45.90 45.60 46.10
DAC 30.12 29.33 27.43 58.06 57.26 58.58 38.20 37.80 37.40 48.50 48.20 48.60
Tamil LIDF 35.01 35.01 35.04 52.87 52.51 54.14 33.20 32.80 32.40 45.80 45.60 46.00
DAC 39.28 38.95 38.47 57.96 58.86 54.45 37.80 37.40 37.00 48.60 48.30 48.70
[…]